Solve for the specified variable.
S =a/1-r for r

Answer by checkley77(12844)

S =(a/1)-r
r=-S+a/1 answer.
OR:
S=a/(1-r)
S(1-r)=a
S-Sr=a
-Sr=a-S
r=(a-S)/-S answer.